# CSGaze: Context-aware Social Gaze Prediction

### v1

@@ 4.3. Implementation Details
We have conducted our experiments on an NVIDIA A100-40GB GPU using the TensorFlow framework. Given the coordinate values of the face, we have used the OpenCV Python library to crop facial regions. The input images, including cropped face images of both individuals, are resized to 224×224. For feature extraction, we employ ResNet34 (resnet34) as the face encoder and ResNet-50 as the scene encoder. KOSMOS-2 (kosmos) served as our MLLM, while RoBERTa (roberta) is used as the text encoder to provide textual embeddings. For the LAEO datasets, since both participants contribute equally to the interaction, equal attention weights (0.5 each) are assigned to the principal ($\alpha_{\text{Principal}}$) and associate ($\alpha_{\text{Associate}}$) participants. The training process is divided into two phases: in Phase-1, the model is pretrained for 10 epochs. During Phase-2 training, we have applied categorical cross-entropy loss for multiclass classification and binary cross-entropy loss for binary classification. The network is optimized using the Adam optimizer with a learning rate of 0.001, and the batch size was set to 128. The model is trained for 200 epochs, with early stopping applied, using a patience value of 5.
Figure 5. Qualitative Analysis. Comparison of our proposed approach, CSGaze, with the baseline method (gaze) across four diverse datasets: GP-Static, AVA-LAEO, UCO-LAEO, and VideoCoAtt. In each example, green bounding boxes indicate the principal face, while pink boxes denote the associate. Predictions are color-coded: blue for correct and red for incorrect outputs.
@@ 4.4.1. Overall and Classwise Prediction Results
Table 2. Performance comparison of SOTA methods using Average Precision (AP) metric on the test sets of the UCO-LAEO and AVA-LAEO datasets. (LAEO: Looking at Each Other). Please note that FG’24 (gupta2024exploring) codebase is not publicly available.
@@
LAEO-Net (marin2011here) 79.5 50.6
AAAI’21 (doosti2021boosting) 65.1 72.2
GP-Static (gaze) 80.3 82.5
FG’24 (gupta2024unified) 95.7 -
MGTR (mgtr) 64.8 66.2
MGTR + GTD-LLM (gtd) 68.3 69.5
LAEO-Net + GTD-LLM (gtd) 83.0 60.4
@@ 4.4.2. Comparison with State-of-the-Art Approaches for the LAEO Class
To evaluate the effectiveness of our proposed approach, CSGaze, across diverse scenarios, we have tested it on the LAEO class using two benchmark datasets: UCO-LAEO and AVA-LAEO. Table 2 compares CSGaze with state-of-the-art methods. CSGaze outperforms all baselines on the AVA-LAEO dataset, demonstrating the effectiveness of contextual information for social gaze prediction. On UCO-LAEO, it performs slightly below the recent method by Gupta et al. (gupta2024unified), likely due to their explicit modeling of gaze interactions and use of multimodal cues like speaking status. Notably, the CSGaze model is lightweight, with approximately 54M parameters, compared to (gupta2024unified) (¿100M). Compared to the baseline (gaze) ( 51M), CSGaze requires slightly more parameters due to the inclusion of cross-attention and contextual modules. Despite this, CSGaze surpasses other methods in multiple aspects, confirming its robustness and generalizability. We believe it is also suitable for real-time deployment.
@@
ChongS (chong2020detecting) 0.78 0.56 0.29
ChongT (chong2020detecting) 0.76 0.53 0.33
Gupta (gupta2022modular) 0.78 0.59 0.36
Baseline (MTGS) (gupta2025mtgs) 0.81 0.60 0.58
@@ 4.4.4. Qualitative Analysis
Figure 5 presents a performance comparison between our proposed approach CSGaze and the baseline method (gaze) across four diverse datasets: GP-Static, AVA-LAEO, UCO-LAEO, and VideoCoAtt. The samples shown are representative examples chosen to illustrate the model’s comparative performance. The results demonstrate that our approach effectively captures social gaze patterns across various scenarios.
@@ 4.4.5. Ablation Study
Comparison of results across various class combinations and reductions: To better understand class-specific performance, we evaluated various class combinations on the GP-Static and VSGaze datasets. Results in Tables 4 and 6 yield the following insights. From Table 4, performance declines as the number of classes increases, indicating greater difficulty in distinguishing between them. Notably, Mutual is consistently the easiest to recognize. In Table 6, Shared Attention (SA) emerges as the most easily classified class, likely due to the presence of group settings and contextual cues such as objects and scene interactions. In contrast, LAEO shows inconsistent performance, likely due to fewer samples. The most challenging distinction is between LAEO and LAH, with the lowest F1-score (0.38). Overall, classification accuracy depends heavily on class pairing—while some (e.g., SA vs. LAEO) are easier to distinguish, others (e.g., LAH vs. LAEO) cause more confusion. These findings emphasize that although more classes increase model complexity, multiclass classification offers better generalizability for real-world scenarios than single-class models.
@@ 4.5. Initial Explainability and Discussion
Our observations indicate that facial features, which incorporate head and eye information, are more predictive for the mutual, single, and miss gaze classes. This is likely because the gaze target in these cases is always another person. In contrast, scene features contribute more significantly to the share and void classes. Share gaze poses a greater challenge, as the gaze targets of both participants can be either another person or an object. This requires scene-object interaction, making the features of the scene more relevant - aligning with the findings of (gupta2025mtgs). Thus, incorporating additional contextual information enhances the accuracy of social gaze prediction.

### v2

@@ 4.3. Implementation Details
We have conducted our experiments on an NVIDIA A100-40GB GPU using the TensorFlow framework. Given the coordinate values of the face, we have used the OpenCV Python library to crop facial regions. The input images, including cropped face images of both individuals, are resized to 224×224. For feature extraction, we employ ResNet34 (He et al., 2016) as the face encoder and ResNet-50 as the scene encoder. KOSMOS-2 (Peng et al., 2023) served as our MLLM, while RoBERTa (Yinhan et al., 2019) is used as the text encoder to provide textual embeddings. For the LAEO datasets, since both participants contribute equally to the interaction, equal attention weights (0.5 each) are assigned to the principal ($\alpha_{\text{Principal}}$) and associate ($\alpha_{\text{Associate}}$) participants. The training process is divided into two phases: in Phase-1, the model is pretrained for 10 epochs. During Phase-2 training, we have applied categorical cross-entropy loss for multiclass classification and binary cross-entropy loss for binary classification. The network is optimized using the Adam optimizer with a learning rate of 0.001, and the batch size was set to 128. The model is trained for 200 epochs, with early stopping applied, using a patience value of 5.
Figure 5. Qualitative Analysis. Comparison of our proposed approach, CSGaze, with the baseline method (Chang et al., 2023) across four diverse datasets: GP-Static, AVA-LAEO, UCO-LAEO, and VideoCoAtt. In each example, green bounding boxes indicate the principal face, while pink boxes denote the associate. Predictions are color-coded: blue for correct and red for incorrect outputs.
@@ 4.4.1. Overall and Classwise Prediction Results
Table 2. Performance comparison of SOTA methods using Average Precision (AP) metric on the test sets of the UCO-LAEO and AVA-LAEO datasets. (LAEO: Looking at Each Other). Please note that FG’24 (Gupta et al., 2024b) codebase is not publicly available.
@@
LAEO-Net (Marin-Jimenez et al., 2011) 79.5 50.6
AAAI’21 (Doosti et al., 2021) 65.1 72.2
GP-Static (Chang et al., 2023) 80.3 82.5
FG’24 (Gupta et al., 2024a) 95.7 -
MGTR (Guo et al., 2022) 64.8 66.2
MGTR + GTD-LLM (Yang and Lu, [n.d.]) 68.3 69.5
LAEO-Net + GTD-LLM (Yang and Lu, [n.d.]) 83.0 60.4
@@ 4.4.2. Comparison with State-of-the-Art Approaches for the LAEO Class
To evaluate the effectiveness of our proposed approach, CSGaze, across diverse scenarios, we have tested it on the LAEO class using two benchmark datasets: UCO-LAEO and AVA-LAEO. Table 2 compares CSGaze with state-of-the-art methods. CSGaze outperforms all baselines on the AVA-LAEO dataset, demonstrating the effectiveness of contextual information for social gaze prediction. On UCO-LAEO, it performs slightly below the recent method by Gupta et al. (Gupta et al., 2024a), likely due to their explicit modeling of gaze interactions and use of multimodal cues like speaking status. Notably, the CSGaze model is lightweight, with approximately 54M parameters, compared to (Gupta et al., 2024a) (¿100M). Compared to the baseline (Chang et al., 2023) ( 51M), CSGaze requires slightly more parameters due to the inclusion of cross-attention and contextual modules. Despite this, CSGaze surpasses other methods in multiple aspects, confirming its robustness and generalizability. We believe it is also suitable for real-time deployment.
@@
ChongS (Chong et al., 2020) 0.78 0.56 0.29
ChongT (Chong et al., 2020) 0.76 0.53 0.33
Gupta (Gupta et al., 2022) 0.78 0.59 0.36
Baseline (MTGS) (Gupta et al., 2025) 0.81 0.60 0.58
@@ 4.4.4. Qualitative Analysis
Figure 5 presents a performance comparison between our proposed approach CSGaze and the baseline method (Chang et al., 2023) across four diverse datasets: GP-Static, AVA-LAEO, UCO-LAEO, and VideoCoAtt. The samples shown are representative examples chosen to illustrate the model’s comparative performance. The results demonstrate that our approach effectively captures social gaze patterns across various scenarios.
@@ 4.4.5. Ablation Study
Comparison of results across various class combinations and reductions. To better understand class-specific performance, we evaluated various class combinations on the GP-Static and VSGaze datasets. Results in Tables 4 and 6 yield the following insights. From Table 4, performance declines as the number of classes increases, indicating greater difficulty in distinguishing between them. Notably, Mutual is consistently the easiest to recognize. In Table 6, Shared Attention (SA) emerges as the most easily classified class, likely due to the presence of group settings and contextual cues such as objects and scene interactions. In contrast, LAEO shows inconsistent performance, likely due to fewer samples. The most challenging distinction is between LAEO and LAH, with the lowest F1-score (0.38). Overall, classification accuracy depends heavily on class pairing—while some (e.g., SA vs. LAEO) are easier to distinguish, others (e.g., LAH vs. LAEO) cause more confusion. These findings emphasize that although more classes increase model complexity, multiclass classification offers better generalizability for real-world scenarios than single-class models.
@@ 4.5. Initial Explainability and Discussion
Our observations indicate that facial features, which incorporate head and eye information, are more predictive for the mutual, single, and miss gaze classes. This is likely because the gaze target in these cases is always another person. In contrast, scene features contribute more significantly to the share and void classes. Share gaze poses a greater challenge, as the gaze targets of both participants can be either another person or an object. This requires scene-object interaction, making the features of the scene more relevant - aligning with the findings of (Gupta et al., 2025). Thus, incorporating additional contextual information enhances the accuracy of social gaze prediction.